\begin{thebibliography}{99}

%\bibitem{ali79}N.D. Alikakos, {\it $L^p$ bounds of solutions of reaction-diffusion equations}, Comm. Partial Differential Equations {\bf 4} (1979) 827-868 

%\bibitem{bl}D. Bartolucchi and C.-S. Lin, {\it Existence and uniqueness for mean field equation on multiply connected domains at the critical parameter}, preprint. 

%\bibitem{bp98}S. Baraket and F. Pacard, {\it Construction of singular limits of a semilinear elliptic equation in dimension $2$}, Calc. Var. PDE {\bf 6} (1998) 1-38. 

\bibitem{bil98}P. Biler, {\it Local and global solvability of some systems modelling chemotaxis}, Adv. Math. Sci. Appl. {\bf 8} (1998) 715-743. 

%\bibitem{bhn94}P. Biler, W. Hebisch, and T. Nadzieja, {\it The Debye system: existence and large time behavior of solutions}, Nonlinear Analysis {\bf 23} (1994) 1189-1209. 

%\bibitem{bhn95}P. Biler, D. Hilhorst, and T. Nadzieja, {\it Existence and nonexistence of solutions for a model of gravitating interaction of particles, II}, Colloq. Math. {\bf 67} (1994) 297-308 

%\bibitem{bkln06a}P. Biler, G. Karch, P. Lauren\c{c}ot, and T. Nadzieja, {\it The $8\pi$-problem for radially symmetric solutions of a chemotaxis model in a disc}, Top. Meth. Nonl. Anal. {\bf 27} (2006) 133-144. 

%\bibitem{bkln06b}P. Biler, G. Karch, P. Lauren\c{c}ot, and T. Nadzieja, {\it The $8\pi$-problem for radially symmetric solutions of a chemotaxis model in the plane}, Math. Meth. Appl. Sci. {\bf 29} (2006) 1563-1583. 

%\bibitem{bcm08}A. Blanchet, J.A. Carrillo, and N. Masmoudi, {\it Infinite time aggregation for the critical two-dimensional Patlak-Keller-Segel model}, Comm. Pure Appl. Math. {\bf 61} (2008) 1449-1481.  

%\bibitem{bdp06}A. Blanchet, J. Dolbeaut, and B. Perthame, {\it Two-dimensional Keller-Segel model: optimal critical mass and qualitative properties of the solutions}, Electronic J. Differential Equations {\bf 2006-44} (2006) 1-33. 

%\bibitem{bs73}H. Brezis and W. Strauss, {\it Semi-linear second-order elliptic equations in $L^1$}, J. Math. Soc. Japan {\bf 25} (1973) 565-590. 

%\bibitem{cf85}L.A. Caffarelli and A. Friedman, {\it Convexity of solutions of semilinear elliptic equations}, Duke Math. J. {\bf 52} (1985) 431-456 

%\bibitem{clmp92}E. Caglioti, P.-L. Lions, C. Marchioro, and M. Pulvirenti, {\it A special class of stationary flows for two-dimensional Euler equations: a statistical mechanics description}, Comm. Math. Phys. {\bf 143} (1992) 501-525. 

%\bibitem{cpy10}D. Cao, S. Peng, and S. Yan, {\it Multiplicity of solutions for the plasma problem in two dimensions}, Adv. Math. {\bf 225} (2010) 2741-2785.  

%\bibitem{ccl03}S.-Y.-A. Chang, C.-C. Chen, and C.-S., Lin, {\it Extremal functions for mean field equation in two dimension}, In; Lectures on Partial Differential Equations: Proceedings in Honor of Luis Nirenbergfs 75th Birthday (ed. S.-Y. A. Chang, C.-S. Lin, and S.-T. Yau), International Press, New York, 2003. 

%\bibitem{cha08}P.-H. Chavanis, {\it Two-dimensional Brownian vortices}, Physica A {\bf 387} (2008) 6917-6942 

%\bibitem{cl91}W. Chen and C. Li, {\it Classification of solutions of some nonlinear elliptic equations}, Duke Math. J. {\bf 63} (1991) 615-622. 

\bibitem{cp81}S. Childress and J.K. Percus, {\it Nonlinear aspects of chemotaxis}, Math. Biosci. {\bf 56} (1981) 217-237. 

%\bibitem{CE}C. Conca and E.E. Espejo, {\it Threshold condition for global existence and blow-up to a radially symmetric drift-diffusion system}, Applied Math Letters (2012) 352-356 

%\bibitem{eks}E.E. Espejo, M. Kurokiba, and T. Suzuki, {\it Blowup threshold and collapse mass separation for a drift-diffusion system in space-dimension two}, preprint 

\bibitem{ds09}J. Dolbeault and C. Schmeiser, {\it The two-dimensional Keller-Segel model after blowup}, Discrete and Contin. Dyn. Syst. Ser. A {\bf 25} (2009) 109-121. 

%\bibitem{ess12}E.E. Espejo, A. Stevens, and T. Suzuki, {\it Simultaneous blowup and mass separation during collapse in an interacting system of chemotactic species}, Differential Integral Equations {\bf 25} (2012) 251-288. 

%\bibitem{esv09}E.E. Espejo, A. Stevens, J.J.L. Vel\'{a}zquez, {\it Simultaneous finite time blow-up in a two-species model for chemotaxis}, Analysis {\bf 29} (2009) 317-338 

%\bibitem{esv10}E.E. Espejo, A. Stevens, J.J.L. Vel\'{a}zques, {\it A note on non-simultaneous blow-up for a drift-diffusion model},  Differential and Integral Equations, {\bf 23} (2010) 451-462 

%\bibitem{egp05}P. Esposito, M. Grossi, and A. Pistoia, {\it On the existence of blowing-up solutions for a mean field equation}, Ann. Inst. H. Poincar\'e Anal. Non Lin\'eaire {\bf 22} (2005) 227-257. 

%\bibitem{Evans} L. EVANS. Partial differential equations.Graduate Studies in Mathematicsm AMS (1998).

%\bibitem{fmp02}M. Fila, H. Matano, and P. Pol\'a\v cik, {\it Existence of $L^1$-connections between equilibria of a semilinear parabolic equation}, J. Dynamical Differential Equations {\bf 19} (2002) 463-491. 

\bibitem{gz98}H. Gajewski and K. Zacharias, {\it Global behaviour of a reaction-diffusion system modelling chemotaxis}, Math. Nachr. {\bf 195} (1998) 77-114.  

%\bibitem{gt}D. Gilbarg and N.S. Trudinger, {\it Elliptic Partial Differential Equations of Second Order}, Springer-Verlag, Berlin, 1983 

%\bibitem{ggos}F. Gladiali, M. Grossi, H. Ohtsuka, and T. Suzuki, {\it Morse indices of multiple blow-up solutions to the two-dimensional Gel'fand problem}, preprint. 

%\bibitem{gos11}M. Grossi, H. Ohtsuka, and T. Suzuki, {\it Asymptotic non-degeneracy of the multiple blow-up solutions to the Gel'fand problem in two space dimensions}, Adv. Differential Equations {\bf 16} (2011) 145-164. 

%\bibitem{gt10}M. Grossi and F. Takahashi, {\it Nonexistence of multi-bubble solutions to some elliptic equations on convex domains}, J. Funct. Anal. {\bf 259} (2010) 904-917. 

\bibitem{hv96}M.A. Herrero and J.J.L. Vel\'azquez, {\it Singularity patterns in a chemotaxis model}, Math. Ann. {\bf 306} (1996) 583-623. 

%\bibitem{henry81}D. Henry, {\it Geometric Theory of Semilinear Parabolic Equations}, Lecture Notes in Math. {\bf 840} Sprinver-Verlag, Berlin, 1981.

%\bibitem{iv99}T. Iwaniec and A. Verde, {\it On the operator ${\cal L}(f) = f \log \vert f\vert$}, J. Funct. Anal. 169 (1999) 391-420 

\bibitem{jl92}W. J\"{a}ger and S. Luckhaus, {\it On explosions of solutions to a system of partial differential equations modelling chemotaxis}, Trans. Amer. Math. Soc. {\bf 329} (1992) 819-824. 

%\bibitem{jm73}G. Joyce and D. Montgomery, {\it Negative temperature states for two-dimensional guiding-center plasma}, J. Plasma Phys. {\bf 10} (1973) 107-121. 

%\bibitem{ks09}N.I. Kavallaris and P. Souplet, {\it Grow-up rate and asymptotics for a two-dimensional Patlak-Keller-Segel model in a disk}, SIAM J. Math. Anal. {\bf 41} (2009) 128-157 

%\bibitem{ks09}N.I. Kavallaris and P. Souplet, {\it Grow-up rate and refined asymptotics for a two-dimensional Patlak-Keller-Segerl model in a disk}, SIAM J. Math. Anal. {\bf 41} (2009) 128-157.  

%\bibitem{ks70}E.F. Keller and L.A. Segel, {\it Initiation of slime mold aggregation viewed as an instability}, J. Theor. Biol.{\bf 26} (1970) 399-415. 

%\bibitem{kie93}M.K.H. Kiessling, {\it Statistical mechanics of classical particles with logarithmic interaction}, Comm. Pure Appl. Math. {\bf 46} (1993) 27-56. 

%\bibitem{ko03}M. Kurokiba and T. Ogawa, {\it Finite time blow-up of the solution for a nonlinear parabolic equation of drift-diffusion type}, Differential and Integral Equations {\bf 4} (2003) 427-452. 

%\bibitem{ko08}M. Kurokiba and T. Ogawa, {\it Wellposedness of the drit-diffusion system in $L^{p}$ arising from the semiconductor device simulation},  J. Math. Anal. Appl. {\bf 342} (2008) 1052-1067 

%\bibitem{kno06}M. Kurokiba, T. Nagai, and T. Ogawa, {\it The uniform boundedness and threshold for the global existence of the radial solution to a drift-diffusion system}, Comm. Pure Appl. Anal. {\bf 5}  (2006) 97-106 

\bibitem{lions84}P.L. Lions, {\it The concentration-compactness principle in the calculus of variation. The locally compact case, Part I}, Ann. Inst. H. Poincar\'e, Analyse nonlin\'eaire {\bf 1} (1984) 109-145. 

%\bibitem{loj63}S. Lojasiewicz, {\it Une propri\`et\'e topologique des sous ensembles analytiques r\'eels}, Coloques du CNRS, Les \'Equations aux D\'riv\'ees Partielles (1963) 117. 

\bibitem{lsv12}S. Luckhaus, Y. Sugiyama, and J.J.L. Vel\'azquez, {\it Measure valued solutions of the 2D Keller-Segel system}, Arch. Rational Mech. Anal. {\bf 206} (2012) 31-80. 

%\bibitem{ms97}N. Mizoguchi and T. Suzuki, {\it Equations of gas combustion: $S$-shaped bifurcation and mushrooms}, J. Differential Equations {\bf 134} (1997) 183-215.  

\bibitem{nag95}T. Nagai, {\it Blow-up of radially symmetric solutions to a chemotaxis system},  Adv. Math. Sci. Appl. {\bf 5} (1995) 581-601. 

\bibitem{nag01}T. Nagai, {\it Blowup of nonradial solutions to parabolic-elliptic systems modeling chemotaxis in two-dimensional domains}, J. Inequal. Appl. {\bf 6} (2001) 37-55. 

%\bibitem{nag11}T. Nagai, {\it Convergence to self-similar solutions for a parabolic-elliptic system of drift-diffusion type in ${\bf R}^2$}, Adv. Differential Equations {\bf 16} (2011) 839-866. 

\bibitem{nsy97}T. Nagai, T. Senba, and K. Yoshida, {\it Application of the Trudinger-Moser inequality to a parabolic system of chemotaxis}, Funkcial. Ekvac. {\bf 40} (1997) 411-433. 

%\bibitem{ns90}K. Nagasaki and T. Suzuki, {\it Asymptotic analysis for two-dimensional elliptic eigenvalues problems with exponentially dominated nonlinearities}, Asymptotic Analysis {\bf 3} (1990) 173-188. 

\bibitem{ns08}Y. Naito and T. Suzuki, {\it Self-similarity in chemotaxis systems}, Colloquium Mathematics {\bf 111} (2008) 11-34. 

%\bibitem{nan73}V. Nanjundiah, {\it Chemotaxis, signal relaying, and aggregation  morphology}, J. Theor. Biol. {\bf 42} (1973) 63-105. 

%\bibitem{os03}H. Ohtsuka and T. Suzuki, {\it Palais-Smale sequence relative to the Trudinger-Moser inequality}, Calc. Vari. in PDE {\bf 17} (2003) 235-255. 

%\bibitem{oss07}H. Ohtsuka, T. Senba, and T. Suzuki, {\it Blowup in infinite time in the simplified system of chemotaxis}, Adv. Math. Sci. Appl. {\bf 17} (2007) 445-472. 

%\bibitem{ons49}L. Onsager, {\it Statistical hydrodynamics}, Suppl. Nuovo Cimento {\bf 6} (1949) 279-287. 

%\bibitem{pl76}Y.B. Pointin and T.S. Lundgren, {\it Statistical mechanics of two-dimensional vortices in a bounded container}, Phys. Fluids (1976) {\bf 19} 1459--1470. 

%\bibitem{rr99}M.M. Rao and Z.D. Ren, {\it Theory of Orlicz Spaces}, Marcel Dekker, New York, 1991.

\bibitem{ssv13}Y. Seki, Y. Sugiyama, and J.J.L. Vel\'azquez, {\it Multiple peak aggregations for the Keller-Segel system}, Nonlinearity {\bf 26} (2013) 319-352. 

\bibitem{sen07}T. Senba, {\it Type II blowup of solutions to a simplified Keller-Segel system in two dimensions}, Nonlinear Analysis {\bf 66} (2007) 1817-1839. 

\bibitem{ss00}T. Senba and T. Suzuki, {\it Some structures of the solution set for a stationary system of chemotaxis}, Adv. Math. Sci. Appl. {\bf 10} (2000) 191-224. 

\bibitem{ss01a}T. Senba and T. Suzuki, {\it Chemotactic collapse in a parabolic-elliptic system of mathematical biology}, Adv. Differential Equations {\bf 6} (2001) 21-50. 

\bibitem{ss01b}T. Senba and T. Suzuki , {\it Parabolic system of chemotaxis: blowup in a finite and the infinite time}, Meth. Appl. Anal {\bf 8} (2001) 349-368.  
\bibitem{ss02a}T. Senba and T. Suzuki, {\it Weak solutions to a parabolic-elliptic system of chemotaxis}, J. Funct. Anal. {\bf 191} (2002) 17-51. 

\bibitem{ss02b}T. Senba and T. Suzuki, {\it Time global solutions to a parabolic-elliptic system modelling chemotaxis}, Asymptotic Analysis {\bf 32} (2002) 63-89. 

%\bibitem{ss03}T. Senba and T. Suzuki, {\it Blowup behavior of solutions to re-scaled J\"ager-Luckhaus system}, Adv. Differential Equations {\bf 8} (2003) 787-820.  

%\bibitem{sim83}L. Simon, {\it Asymptotics for a class of nonlinear evolution equations with applications to geometric problems}, Ann. Math {\bf 118} (1983) 525-571. 

\bibitem{sc02}C. Sire and P.-H. Chavanis, {\it Thermodynamics and collapse of self-gravitating Brownian particles in $D$ dimensions}, Phys. Rev. E {\bf 66} (2002) 046133. 

%\bibitem{sw05}I. Shafrir, G. Wolansky, {\it Moser-Trudinger and logarithmic HLS inequalities for systems}, J. Euro. Math. Soc. {\bf 7} (2005) 413-448 

\bibitem{struwe}M. Struwe, {\it Variational Methods, Applications to Nonlinear Partial Differential Equations and Hamilton Systems}, third edition, Springer Verlag, Berlin, 2000. 

%\bibitem{s92}T. Suzuki, {\it Global analysis for a two-dimensional elliptic eigenvalue problem with exponential nonlinearity}, Ann. Inst. H. Poincar\'e Anal. Non Lin\'eaire {\bf 9} (1992) 367-398. 

%\bibitem{suzuki94}T. Suzuki, {\it Semilinear Elliptic Equations}, Gakkotosho, Tokyo, 1994.  

\bibitem{suzuki05}T. Suzuki, {\it Free Energy and Self-Interacting Particles}, Birkh\"auser, Boston, 2005. 

%\bibitem{suzuki08}T. Suzuki, {\it Mean Field Theories and Dual Variation}, Atlantis Press, Amsterdam-Paris, 2008. 

%\bibitem{s}T. Suzuki, {\it 2D Brownian point vortices  and the drift-diffusion model}, preprint

\bibitem{s13}T. Suzuki, {\it Exclusion of boundary blowup for 2D chemotaxis system provided with the Dirichlet boundary condition for the Poisson part}, J. Math. Pure Appl. 100 (2013) 347-367. 

\bibitem{s14}T. Suzuki, {\it Almost collapse mass quantization in 2D Smoluchowski-Poisson equation}, arXive1311.5679. 

%\bibitem{sn88}T. Suzuki and K. Nagasaki, {\it On the nonlinear eigenvalue problem $\Delta u+\lambda e^u=0$}, Trans. Amer. Math. Soc. {\bf 39} (1988) 591-608.  

\bibitem{suzuki11}T. Suzuki and T. Senba, {\it Applied Analysis - Mathematical Methods in Natural Science}, 2nd edition, Imperial College Press, 2011. 

%\bibitem{sy}T. Suzuki and Y. Yamada, {\it Global-in-time behavior of Lotka-Volterra system with diffusion}, preprint. 

%\bibitem{win}M. Winkler, {\it Finite-time blow-up in the higher-dimensional parabolic-parabolic Keller-Segel system}, preprint. 

%\bibitem{wol02}G. Wolansky, {\it Multi-components chemotactic system in the absence of conflicts},  Euro. J. Appl. Math. {\bf 3} (2002) 641-661 

\end{thebibliography}

\documentclass[11pt]{article}

\usepackage[mathscr]{eucal}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsfonts}
\usepackage{amscd}
\usepackage{amsthm}
\usepackage{latexsym}
\usepackage{graphicx}

\def\be{\begin{equation}}
\def\ee{\end{equation}}
\def\bse{\begin{subequations}}
\def\ese{\end{subequations}}

\newtheorem{thm}{Theorem}

\newtheorem{lem}{Lemma}[section]
\newtheorem{prop}{Proposition}[section]

\def\bse{\begin{subequations}}
\def\ese{\end{subequations}}

\usepackage{geometry}

\def\XXint#1#2#3{{
\setbox0=\hbox{$#1{#2#3}{\int}$}
\vcenter{\hbox{$#2#3$}}\kern-.5\wd0}}
\title{Residual vanishing for blowup solutions to 2D Smoluchowski-Poisson equation}

\author{Takashi Suzuki} 

\date{\today}

\begin{document}
\let\cleardoublepage\clearpage%svinei kenes selides

\maketitle

%\begin{center} 
%{\it dedicated to Professor Toshitaka Nagai}
%\end{center} 

\begin{abstract}
We study Smoluchowski-Poisson equation in two space dimensions provided with Dirichlet boundary condition for the Poisson part.  For this equation several profiles of blowup solution have been noticed.  Here we show the residual vanishing. 
\end{abstract}

\section{Introduction}\label{sec:1}

We study parabolic-elliptic system  proposed in statistical physics to describe the motion of mean field of many self-gravitating Brownian particles \cite{sc02}.  It is composed of the Smoluchowski part 
\begin{equation}
u_t=\Delta u-\nabla \cdot u\nabla v \quad \mbox{in $\Omega\times(0,T)$} 
 \label{s}
\end{equation} 
with null-flux boundary condition 
\begin{equation}
\frac{\partial u}{\partial \nu}-u\frac{\partial v}{\partial\nu}=0 \quad \mbox{on $\partial\Omega \times (0,T)$} 
 \label{eqn:s3} 
\end{equation}
and the Poisson part in the form of 
\begin{equation}
-\Delta v=u, \quad \left. v\right\vert_{\partial\Omega}=0, 
 \label{eqn:s4} 
\end{equation}
where $\Omega\subset {\bf R}^2$ is a bounded domain with smooth boundary $\partial\Omega$ and $\nu$ is the outer unit normal vector.  Initial condition is given as 
\begin{equation} 
\left. u\right\vert_{t=0}=u_{0}(x)\geq 0 \quad \mbox{in $\Omega$}, 
 \label{eqn:s2}
\end{equation} 
where $u_0=u_0(x)$ is a smooth function.  

System (\ref{s})-(\ref{eqn:s2}) is subject to thermodynamical laws, {\it total mass conservation} and {\it free energy decreasing}, 
\begin{eqnarray} 
& & \frac{d}{dt}\int_\Omega u=\int_\Omega \nabla\cdot (\nabla u-u\nabla v)=\int_{\partial\Omega}\frac{\partial u}{\partial \nu}-u\frac{\partial v}{\partial\nu} \ ds=0 
 \label{eqn:tmc} \\ 
& & \frac{d}{dt}{\cal F}(u)=-\int_\Omega u\vert \nabla (\log u-v)\vert^2\leq 0  
 \label{eqn:fed}
\end{eqnarray} 
where $ds$ denotes the surface element and 
\begin{equation} 
{\cal F}(u)=\int_\Omega u(\log u-1) \ dx -\frac{1}{2}\langle (-\Delta)^{-1}u,u\rangle 
 \label{eqn:free-e}
\end{equation} 
with $v=(-\Delta)^{-1}u$ standing for (\ref{eqn:s4}).  

A related model arises in the context of chemotaxis in theoretical biology \cite{jl92, nag95}, where the Poisson part is provided with the Neumann boundary condition such as 
\begin{equation}
-\Delta v=u-\frac{1}{\vert\Omega\vert}\int_\Omega u, \quad \left. \frac{\partial v}{\partial \nu}\right\vert_{\partial\Omega}=0, \quad \int_\Omega v=0. 
 \label{eqn:sd} 
\end{equation}
Concerning (\ref{s})-(\ref{eqn:s3}), (\ref{eqn:s2}), and (\ref{eqn:sd}), there is a threshold of $\Vert u_0\Vert_1=\lambda$ for the blowup of the solution.  More precisely, if $\lambda<4\pi$ the solution exists global-in-time \cite{bil98, gz98, nsy97}.  If a local mass greater than $4\pi$ is concentrated on a boundary point, on the contrary, there arises blowup in finite time \cite{nag01, ss01b}. Underlying blowup mechanisms were suspected from the study of stationary solutions \cite{cp81}. This attempt was followed by \cite{hv96, ss00}, using radially symmetric and general stationary solutions, respectively.  Up to now several properties have been known \cite{ss01a, sen07, ns08}.  See our previous work \cite{s14} and the references therein.  System (\ref{s})-(\ref{eqn:s2}), provided with Dirichlet condition for the Poisson part, is taken by \cite{s13}. It excludes boundary blowup points. Here we continue the study \cite{s14} on interior blowup points.  

Fundamental features of system (\ref{s})-(\ref{eqn:s2}) are the following. First, local-in-time unique existence of the classical solution is standard, given smooth initial value $u_0=u_0(x)\geq 0$.  Henceforth, $T\in (0, +\infty]$ denotes its maximal existence time. If $u_0\not\equiv 0$, which we always assume, the strong maximum principle and the Hopf lemma guarantee $u(\cdot,t)>0$ on $\overline{\Omega}$ for $t>0$.  Maximal existence time $T$ of non-stationary solution $u=u(\cdot,t)$, on the other hand, is estimated from below by $\Vert u_0\Vert_\infty$.  Hence $T<+\infty$ implies 
\[ \lim_{t\uparrow T}\Vert u(\cdot,t)\Vert_\infty=+\infty \] 
and the blowup set 
\begin{equation} 
{\cal S}=\{ x_0\in\overline{\Omega} \mid \mbox{$\exists x_k\rightarrow x_0$, $\exists t_k\uparrow T$ such that $u(x_k,t_k)\rightarrow+\infty$}\}  
 \label{eqn:blowupset1}
\end{equation} 
is non-empty.  Since boundary blowup points are excluded \cite{s13}, if $T<+\infty$ in (\ref{s})-(\ref{eqn:s2}) we have 
\begin{equation} 
u(x,t)dx\rightharpoonup \sum_{x_0\in {\cal S}}m(x_0)\delta_{x_0}(dx)+f(x)dx \quad \mbox{in ${\cal M}(\overline{\Omega})=C(\overline{\Omega})'$} 
 \label{eqn:bfq}
\end{equation} 
as $t\uparrow T$.  Here, the blowup set satisfies ${\cal S}\subset \Omega$ with $\sharp {\cal S}<+\infty$, and it holds that $0\leq f=f(x)\in L^1(\Omega)\cap C(\overline{\Omega}\setminus {\cal S})$.  

The blowup mechanism at each inner blowup point, described by \cite{s14}, is more complicated than the ones suspected before. Let $x_0\in{\cal S}$ and $R(t)=(T-t)^{1/2}$. As is shown in \cite{suzuki05, suzuki11} it holds that 
\begin{equation} 
\lim_{b\uparrow+\infty}\limsup_{t\uparrow T}\left\vert \Vert u(\cdot,t)\Vert_{L^1(B(x_0, bR(t))}-m(x_0)\right\vert =0. 
 \label{eqn:ext8}
\end{equation} 
Henceforth, $C_i$, $i=1,2,\cdots, 15$, denote positive constants. 

\begin{thm}[\cite{s14}]
Any $t_k\uparrow T$ admits a sub-sequence, denoted by the same symbol, and $m\in {\bf N}\cup \{0\}$, provided with the following property. Thus, given $0<\varepsilon\ll 1$ and $R\gg 1$, there is $\tilde s\geq 1$ such that, for $t_k'\uparrow T$ defined by $R(t_k')=\tilde sR(t_k)$ we have $0<r^j_k\leq C_1R(t_k')$ and $x^j_k\in B(x_0, C_1R(t_k'))$, $1\leq j\leq m$.  Then, for $B_k^j=B(x_k^j, b_j)$ and $\displaystyle{ G_k=\bigcup_{j=1}^mB_k^j}$, it holds that 
\begin{eqnarray} 
& & B_k^i\cap B_k^j=\emptyset, \quad i\neq j, \ k \gg 1 \nonumber\\ 
& & \limsup_{k\rightarrow\infty}\left\vert \Vert u(\cdot,t_k')\Vert_{L^1(B_k^j)}-8\pi\right\vert <\varepsilon, \quad 1\leq j\leq m \nonumber\\ 
& & \lim_{b\uparrow+\infty}\limsup_{k\rightarrow \infty}R(t_k')^2\Vert u(\cdot,t_k')\Vert_{L^\infty(B(x_0, bR(t_k'))\setminus G_k)}\leq C_2R^{-2}. 
 \label{eqn:1.209}
\end{eqnarray} 
 \label{thm:1.11}
\end{thm}

In this paper we show that if (\ref{eqn:1.209}) is extended continuously then it holds that $m(x_0)\in 8\pi {\bf N}$.  Namely, here we assume that any $\varepsilon>0$ admits $m_j(t)\in {\bf N}\cup \{ 0\}$, $x_j(t)\in B(x_0, C_3R(t))$, and $0<r_j(t)\leq C_3R(t)$ for $0<T-t\ll 1$ and $1\leq j\leq m(t)\in {\bf N}\cup \{0\}$, such that, for $B_j(t)=B(x_j(t), r_j(t))$ and $\displaystyle{ G(t)=\bigcup_{j=1}^{m(t)}B_j(t)}$ it holds that 
\begin{eqnarray} 
& & B_i(t)\cap B_j(t)=\emptyset, \quad i\neq j \nonumber\\ 
& & \limsup_{t\uparrow T}\max_{1\leq j\leq m(t)}\left\vert \Vert u(\cdot,t)\Vert_{L^1(B_j(t))}-8\pi\right\vert <\varepsilon \nonumber\\ 
& & \lim_{b\uparrow+\infty}\limsup_{t\uparrow T}R(t)^2\Vert u(\cdot,t)\Vert_{L^\infty(B(x_0, bR(t))\setminus G(t))}\leq C_4.  
 \label{eqn:1.209ih}
\end{eqnarray}  
Then our result arises as follows. 

\begin{thm}
If (\ref{eqn:1.209ih}) holds in (\ref{s})-(\ref{eqn:s2}) with $T<+\infty$, then there is $\tilde t_k\uparrow T$ such that 
\begin{equation} 
\lim_{b\uparrow +\infty}\lim_{k\rightarrow\infty}\Vert u(\cdot,\tilde t_k)\Vert_{L^1(B(x_0, bR(\tilde t_k))\setminus G(\tilde t_k))}<\varepsilon. 
 \label{eqn:2.189ih}
\end{equation} 
Hence we obtain $m(x_0)=8\pi m$, $m\in {\bf N}$, in (\ref{eqn:bfq}), and in particular, $m(t)=m$ in (\ref{eqn:1.209ih}). 
 \label{thm:2.14ih}
\end{thm} 

We call (\ref{eqn:2.189ih}) the {\it residual vanishing}. Although $m(x_0)\in 8\pi {\bf N}$ follows from (\ref{eqn:ext8}), (\ref{eqn:1.209ih}), and (\ref{eqn:2.189ih}), we shall show $m(x_0)\in 8\pi {\bf N}$ first, and then (\ref{eqn:2.189ih}). In future we shall discuss the problems to derive (\ref{eqn:1.209ih}) from (\ref{eqn:1.209}) and also to refine (\ref{eqn:2.189ih}) to 
\[ \displaystyle{ \lim_{b\uparrow+\infty}\limsup_{t\uparrow T}\Vert u(\cdot,t)\Vert_{L^1(B(x_0, bR(t))\setminus G(t))}<\varepsilon}. \] 

This paper is composed of three sections. Taking preliminaries in the next section, we prove Theorem \ref{thm:2.14ih} in the final section. 

\section{Prelimiaries}

Weak solution is a fundamental tool in later arguments. This notion was introduced first for the pre-scaled Smoluchowski-Poisson equation \cite{ss02a}. Let $G=G(x,x')$ be the Green's function to (\ref{eqn:s4}).  Then we say that $0\leq \mu=\mu(dx,t)\in C_\ast([0,T], {\cal M}(\overline{\Omega}))$ is a weak solution to (\ref{s})-(\ref{eqn:s4}) if there is 
\[ 0\leq \nu=\nu(\cdot,t)\in L^\infty_\ast(0,T; {\cal E}') \] 
called multiplicate operator satisfying the following properties, where ${\cal E}$ is the closure of the linear space 
\[ {\cal E}_0=\{ \psi+\rho_\varphi \mid \psi\in C(\overline{\Omega}\times\overline{\Omega}), \ \varphi\in {\cal X} \}, \quad {\cal X}=\{ \varphi\in C^2(\overline{\Omega}) \mid \left. \frac{\partial\varphi}{\partial \nu}\right\vert_{\partial\Omega}=0\} \] 
in $L^\infty(\Omega\times\Omega)$ and $\rho_\varphi(x,x')=\nabla\varphi(x)\cdot \nabla_xG(x,x')+\nabla\varphi(x')\cdot\nabla_{x'}G(x,x')$: 
\begin{enumerate} 
\item For $\varphi\in {\cal X}$ the mapping $t\in [0,T]\mapsto \langle \varphi, \mu(dx,t)\rangle$ is absolutely continuous and there holds 
\begin{equation} 
\frac{d}{dt}\langle \varphi, \mu(dx,t)\rangle=\langle \Delta \varphi, \mu(dx,t) \rangle+\frac{1}{2}\langle \rho_\varphi, \nu(\cdot,t) \rangle_{{\cal E}, {\cal E}'} \quad \mbox{a.e. $t$}. 
 \label{eqn:mul1}
\end{equation} 
\item We have 
\begin{equation} 
\left. \nu(\cdot,t)\right\vert_{C(\overline{\Omega}\times\overline{\Omega})}=\mu(dx,t)\otimes \mu(dx',t) \quad \mbox{a.e. $t$}.  
 \label{eqn:mul2}
\end{equation} 
\end{enumerate} 
Here we confirm that the property $\nu\geq 0$ of $\nu\in {\cal E}'$ means 
\[ \left\vert \langle f, \nu\rangle_{{\cal E}, {\cal E}'}\right\vert\leq \langle g, \nu \rangle \] 
for any $f, g\in {\cal E}$ satisfying $\vert f\vert\leq g$ a.e. in $\Omega\times\Omega$.  

We note the following properties.  First, the total mass conservation of this weak solution  
\[ \mu(\overline{\Omega},t)=\mu(\overline{\Omega}, 0), \quad t\in [0,T]  \] 
is obvious.  Next, this weak solution cannot be a measure-valued solution constructed in \cite{ds09, lsv12, ssv13}.  In fact, any collision of collapses are not admitted here, and more precisely, we have the following property.  
\begin{lem}[\cite{ss02a}] 
If the initial meause $\mu_0(dx)\in {\cal M}(\overline{\Omega})$ admits $x_0\in \Omega$ such that 
\[ \mu_0(\{ x_0\})>8\pi, \quad \lim_{R\downarrow 0}\frac{1}{R^2}\left\langle \vert x-x_0\vert^2\chi_{B(x_0,R)}, \mu_0(dx)\right\rangle =0 \] 
then there is no weak solution to (\ref{s})-(\ref{eqn:s4}) even local-in-time.   \label{thm:ss02a}
\end{lem} 
It is, however, provided with the following property, derived from the fact that ${\cal E}$ is separable.  

\begin{lem}[\cite{ss02a}]
Let $\{ \mu_k(dx,t)\}\subset C_\ast([0,T], {\cal M}(\overline{\Omega}))$ be a sequence of weak solutions to (\ref{s})-(\ref{eqn:s4}). Let the associated multiplicate operator of $\mu_k(dx,t)$ be $\nu_k(\cdot,t)\in L^\infty_\ast(0,T; {\cal E}')$, and assume 
\begin{equation} 
\mu_{k}(\overline{\Omega}, 0)+\sup_{t\in [0,T]}\Vert \nu_k(\cdot,t)\Vert_{{\cal E}'} \leq C_{5}, \quad k=1,2,\cdots. 
 \label{eqn:w-bound}
\end{equation} 
Then we have a subsequence denoted by the same symbol, $\mu(dx,t)\in C_\ast([0,T], {\cal M}(\overline{\Omega}))$, and $\nu(\cdot,t)\in L^\infty_\ast(0, T; {\cal E}')$ such that 
\begin{eqnarray*} 
& & \mu_k(dx,t)\rightharpoonup \mu(dx,t) \quad \mbox{in $C_\ast([0,T], {\cal M}(\overline{\Omega}))$} \\ 
& & \nu_k(\cdot,t)\rightharpoonup \nu(\cdot,t) \quad \qquad \mbox{in $L^\infty_\ast(0,T; {\cal E}')$}  
\end{eqnarray*} 
up to a sub-sequence, and this $\mu(dx,t)$ is a weak solution to (\ref{s})-(\ref{eqn:s4}) with the multiplicate operator $\nu(\cdot,t)$ satisfying 
\[ \mu(\overline{\Omega},0)+\Vert \nu(\cdot,t)\Vert_{{\cal E}'}\leq C_{5}.  \] 
 \label{pro:1}
\end{lem} 

We agree with the following notations. First, if $\mu(dx,t)$ has a density as 
\[ \mu(dx,t)=u(x,t)dx, \quad 0\leq u(\cdot,t)\in C([0,T], L^1(\Omega)), \] 
then the multiplicate operator is always taken as 
\[ \nu(\cdot,t)=u(x,t)u(x',t) \ dxdx', \] 
recalling ${\cal E}\subset L^\infty(\Omega\times\Omega)$.  Under this agreement, condition (\ref{eqn:w-bound}) is reduced to 
\begin{equation} 
\mu_{k}(\overline{\Omega},0)\leq C_{6}, \quad k=1,2,\cdots 
 \label{eqn:sk1} 
\end{equation}
if each $\mu_k(dx,t)$ takes density in $[0,T)$ such as 
\[ \mu_k(dx,t)=u_k(x,t)dx, \quad 0\leq u_k=u_k(\cdot,t)\in C([0,T), L^1(\Omega)). \] 
In fact, since we take 
\[ \nu_k(\cdot,t)=u_k(x,t)u_k(x',t)dxdx' \] 
in this case, inequality (\ref{eqn:sk1}) means 
\[ \Vert u_k(\cdot,t)\Vert_1=\Vert u_{k}(\cdot,0)\Vert_1=\mu_{k}(\overline{\Omega},0)\equiv \lambda_k \leq C_{6}. \] 
Therefore, (\ref{eqn:w-bound}) follows with 
\[ \Vert \nu(\cdot,t)\Vert_{{\cal E}'}=\lambda_k^2\leq C_{6}^2. \] 
Using this property, we shall derive a hierarchy of weak solutions in later arguments.  

We can define also the regularity of the weak solution.  First, given $\mu=\mu(\cdot, t)\in {\cal M}(\overline{\Omega})$, we have a unique $v=v(\cdot,t)\in W^{1,q}(\Omega)$, $1\leq q<2$, such that 
\[ -\Delta v=\mu, \quad \left. v\right\vert_{\partial\Omega}=0. \] 
Let $I\subset (0,T)$ be an open interval and $\omega\subset \Omega$ an open set. If the weak solution $\mu(dx,t)$ has a density $u=u(\cdot,t)\in L^p(\omega)$ in $\omega\subset \Omega$, $1<p<\infty$, for $t\in I$, the above $v=v(\cdot,t)$ is in $W^{2,p}_{loc}(\omega)$ from the elliptic regularity. By Sobolev's and Morrey's imbedding theorems, this regularity implies $(u\nabla v)(\cdot,t)\in L^1_{loc}(\omega)$.  Then we assign 
\[ \frac{d}{dt}\langle \varphi, \mu(dx,t)\rangle=\langle \Delta \varphi(dx,t)\rangle+\langle \nabla \varphi\cdot\nabla v, \mu(dx,t)\rangle, \quad \mbox{a.e. $t\in I$} \] 
for any $\varphi\in C_0^2(\omega)$.  In such a case we say that $\mu(dx,t)$ is regular in $\omega\times I$. In Lemma \ref{pro:1}, if $\mu_k(dx,t)$ is regular with the density $u_k(x,t)$ in $\omega\times(0,T)$ satisfying 
\[ \sup_{t\in [0,T]}\Vert u_k(\cdot,t)\Vert_{L^p(\omega)}\leq C_{7} \] 
for $p>1$, then the generated $\mu(dx,t)$ is also regular in $\omega\times (0,T)$.  Conversely, we have the following properties by the $\varepsilon$ regularity \cite{ss02b, s14}.  First, if the weak solution $\mu(dx,t)\in C_\ast([0,T], {\cal M}(\overline{\Omega}))$ is generated by a sequence of classical solutions $\{u_k(x,t)\}$ then its singular part $\mu_s(dx,t)$ is composed of a finite sum of delta functions.  Furthermore, if $\mu(dx,t_0)$, $0<t_0<T$, is regular in the sense of measure in an open set $\hat \omega\subset \Omega$, then it is regular in the above sense.  More precisely, $\mu(dx,t)$ takes a smooth density function $f=f(x,t)$ in $\omega\times (t_0-\delta, t_0+\delta)$, where $\omega$ is an open set satisfying $\overline{\omega}\subset \hat \omega$ and $0<\delta \ll 1$. 

The weak solution $a(dx,t)\in C_\ast(-\infty,+\infty; {\cal M}({\bf R}^2))$ to 
\begin{equation} 
a_t=\Delta a-\nabla\cdot a\nabla\Gamma\ast a \quad \mbox{in ${\bf R}^2\times (-\infty, +\infty)$} 
 \label{eqn:entire}
\end{equation}
is defined similarly, where $\displaystyle{ \Gamma(x)=\frac{1}{2\pi}\log\frac{1}{\vert x\vert}}$ is the fundamental solution to $-\Delta$, ${\cal M}({\bf R}^2)=C_0({\bf R}^2)'$ with 
\[ C_0({\bf R}^2)=\{ f\in C({\bf R}^2\bigcup\{\infty\}) \mid f(\infty)=0 \}, \] 
and ${\bf R}^2\bigcup \{\infty\}$ denotes one-point compactification of ${\bf R}^2$. Then the following property is shown, see \cite{s14}. 

\begin{prop}[Liouville property]
Let $0\leq a=a(dx,t)\in C_\ast((-\infty,+\infty), {\cal M}({\bf R}^2))$ be a weak solution to (\ref{eqn:entire}) with uniformly bounded multiplicate operator. Then we have either $a({\bf R}^2,t)=8\pi$ or $a({\bf R}^2, t)=0$, exclusively in $t\in {\bf R}$. 
 \label{lem:1.2} 
\end{prop} 

We can also define the weak solution $\zeta(dy,s)\in C_\ast(-\infty,+\infty; {\cal M}({\bf R}^2))$ to 
\begin{equation} 
\zeta_s=\Delta \zeta-\nabla\cdot \zeta\nabla(\Gamma\ast \zeta+\vert y\vert^2/4) \quad \mbox{in ${\bf R}^2\times (-\infty, +\infty)$},   
 \label{eqn:wsl1}
\end{equation} 
which arises as the weak scaling limit of $u=u(x,t)$.  Thus, given $x_0\in{\cal S}\subset \Omega$, we take the backward self-similar transformation 
\[ z(y,s)=(T-t)u(x,t), \quad y=(x-x_0)/(T-t)^{1/2}, \quad s=-\log(T-t). \] 
Let $t_k\uparrow+\infty$ and put 
\[ s_k=-\log (T-t_k) \ \uparrow+\infty. \] 
Then, passing to a sub-sequence denoted by the same symbol, we have 
\begin{equation} 
z(y, s+s_k)dy\rightharpoonup \zeta(dy,s) \quad \mbox{in $C_\ast(-\infty, +\infty; {\cal M}({\bf R}^2)$}), 
 \label{eqn:zeta}
\end{equation} 
where $\zeta(dy,s)$ is weak solution to (\ref{eqn:wsl1}) provided with a uniformly bounded multiplicate operator.  In (\ref{eqn:zeta}), the important property called {\it parabolic envelope} arises as 
\begin{equation} 
\zeta({\bf R}^2, s)=m(x_0)>0, \quad \langle \vert y\vert^2, \zeta(dy,s)\rangle \leq C_{8}
 \label{eqn:pe}
\end{equation} 
valid to $s\in (-\infty, +\infty)$ with $m(x_0)>0$ defined by (\ref{eqn:bfq}), see \cite{suzuki05, s13}. 

Here we take the {\it scaling back} of $\zeta(dy,s)$, defined by the transformation 
\begin{equation} 
A(dy', s')=e^s\zeta(dy,s), \quad y'=e^{-s/2}y, \quad s'=-e^{-s}. 
 \label{eqn:c9}
\end{equation} 
It has an extension as $0\leq A=A(dy, s)\in C_\ast((-\infty, 0], {\cal M}({\bf R}^2))$ with $A(dy,0)=m(x_0)\delta_0(dy)$. It becomes also a weak solution to 
\begin{equation} 
A_s=\Delta A-\nabla\cdot A\nabla\Gamma\ast A \quad \mbox{in ${\bf R}^2\times (-\infty, 0)$} 
 \label{eqn:c7}
\end{equation}  
satisfying 
\begin{equation} 
A({\bf R}^2, s)=m(x_0), \quad -\infty<s<0
 \label{eqn:c8}
\end{equation}  
with a uniformly bounded multiplicate operator. 

Now, given $\tilde s_\ell\uparrow +\infty$, we take 
\begin{equation} 
A_\ell(dy)=A(dy,-\tilde s_\ell)/m(x_0) 
 \label{eqn:52}
\end{equation} 
to apply {\it concentration compactness principle} \cite{lions84} (see also p. 39 of \cite{struwe}).  Then we obtain the following lemma, which implies Theorem \ref{thm:1.11}. 

\begin{lem}[concentration compactness]
Passing to a sub-sequence we have $m\in {\bf N}\cup\{0\}$ such that any $\varepsilon>0$ admits $y_\ell^j\in {\bf R}^2$ and $b_j>0$, $1\leq j\leq m$, satisfying 
\begin{eqnarray} 
& & \lim_{\ell\rightarrow\infty}\vert y^i_\ell-y^j_\ell\vert=+\infty, \quad \forall i\neq j \label{eqn:208-1} \\ 
& & \limsup_{\ell\rightarrow \infty}\vert A_\ell (B_\ell^j))-8\pi\vert <\varepsilon, \quad \forall j \label{eqn:208-2} \\ 
& & \vert y^j_\ell\vert \leq C_{9}(1+\max_j b_j)\tilde s_\ell^{1/2}, \quad \forall \ell\gg 1, \ \forall j \label{eqn:211}
\end{eqnarray} 
for $B_\ell^j=B(y_\ell^j, b_j)$.  Furthermore, there arises one of the following alternatives. 
\begin{enumerate}
\item $m(x_0)>8\pi m+\varepsilon$ and $A_\ell$, $\ell\gg 1$, is regular in $\displaystyle{{\bf R}^2\setminus \bigcup_{j=1}^mB(y_\ell^j, b_j)}$.  It holds that 
\begin{eqnarray} 
& & \liminf_{\ell\rightarrow\infty}A_\ell\left( {\bf R}^2\setminus \bigcup_{\ell=1}^mB_\ell^j\right)\geq m(x_0)-8\pi m-\varepsilon \label{eqn:209-2} \\ 
& & \lim_{\ell\rightarrow\infty}\Vert A_\ell\Vert_{L^\infty({\bf R}^2\setminus \bigcup_{j=1}^mB_\ell^j)}=0.  \label{eqn:209-1}
\end{eqnarray} 
\item $m(x_0)=8\pi m$ and 
\begin{equation} 
\limsup_{\ell\rightarrow\infty}A_\ell\left( {\bf R}^2\setminus \bigcup_{j=1}^mB(y_\ell^j, b_j)\right)<\varepsilon. 
 \label{eqn:210}
\end{equation} 
\end{enumerate} 
 \label{lem:3.1}
\end{lem} 

\section{Proof of Theorem \ref{thm:2.14ih}} 

From the assumption (\ref{eqn:1.209ih}), $\zeta(dy,s)$ generated in the previous section satisfies an additional condition.  Namely, each $0<\varepsilon\ll 1$ admits $s_1\gg 1$ provided with the following properties.  First, for $\tilde s\geq s_1$ there are $m(\tilde s)\in {\bf N}\cup \{0\}$, $y_j(\tilde s)\in {\bf R}^2$, and $b_j(\tilde s)>0$, $1\leq j\leq m(\tilde s)$, such that $\vert y_j(\tilde s)\vert\leq C_{10}\tilde s^{1/2}$ and $b_j(\tilde s)\leq C_{10}$. Next, $\zeta(dy,-\log \tilde s)$ is regular in $\displaystyle{ \bigcup_{\tilde s\geq \tilde s_1}({\bf R}^2\setminus E_{\tilde s})\times \{ -\log \tilde s\}}$ for 
\begin{equation} 
 E_{\tilde s}=\bigcup_{j=1}^{m(\tilde s)}B_j(\tilde s), \quad \tilde B_j(\tilde s)=B(\tilde s^{-1/2}y_j(\tilde s), \tilde s^{-1/2}b_j(\tilde s)). 
 \label{eqn:ext14}
\end{equation} 
Finally, it holds that $B_i(\tilde s)\cap B_j(\tilde s)=\emptyset$, $i\neq j$, and 
\begin{eqnarray} 
& & \sup_{\tilde s\geq s_1}\Vert \zeta(dy, -\log \tilde s)\Vert_{L^\infty({\bf R}^2\setminus E_{\tilde s})}\leq C_{11} \nonumber\\ 
& & \sup_{\tilde s\geq s_1, \ 1\leq j\leq m(\tilde s)}\left\vert \zeta(B_j(\tilde s), -\log \tilde s)-8\pi\right\vert <\varepsilon. 
 \label{eqn:2.200ih}
\end{eqnarray} 

If $m(x_0)\not\in 8\pi {\bf N}$, we have always the first alternative in Lemma \ref{lem:3.1}, which implies the existence of $\delta>0$ such that 
\begin{equation} 
\inf_{s\geq \tilde s_1}\zeta ({\bf R}^2\setminus E_{\tilde s}, -\log \tilde s)\geq \delta. 
 \label{eqn:ext13}
\end{equation} 
If (\ref{eqn:ext13}) is not the case there is $s_2\geq s_1$ such that 
\[ \zeta ({\bf R}^2\setminus E_{s_2}, -\log s_2) <\varepsilon. \] 
Then it holds that (\ref{eqn:2.189ih}) with some $\tilde t_k\uparrow T$.  Hence Theorem \ref{thm:2.14ih} is reduced to the following lemma.  For the proof we use (\ref{eqn:wsl1}), particularly, the term $\vert y\vert^2/4$, which attract the density of $\zeta(dy,s)$ to $\vert y\vert=\infty$.  

\begin{lem}
The weak scaling limit $\zeta(dy,s)\in C_\ast(-\infty, +\infty; {\cal M}({\bf R}^2))$, generated by (\ref{eqn:zeta}), does not satisfy (\ref{eqn:2.200ih}) and (\ref{eqn:ext13}), simultaneously. 
 \label{lem:2.10ih}
\end{lem}

\begin{proof} 
Similarly to Lemma \ref{pro:1} concerning (\ref{s})-(\ref{eqn:s4}), given $\tilde s_\ell\uparrow+\infty$, we have a sub-sequence denoted by the same symbol such that 
\begin{equation} 
\zeta(dy,s-\tilde s_\ell)\rightharpoonup \tilde \zeta(dy,s) \quad \mbox{in $C_\ast(-\infty, +\infty; {\cal M}({\bf R}^2))$}. 
 \label{eqn:2.204ih}
\end{equation} 
This $\tilde \zeta(dy,s)$ is a weak solution to (\ref{eqn:wsl1}). Furthermore, since $\{ \zeta(dy,s-\tilde s_\ell)\}$ is tight by (\ref{eqn:pe}), it holds that 
\begin{equation} 
\tilde \zeta({\bf R}^2,s)=m(x_0), \quad \langle \vert y\vert^2, \tilde \zeta(dy,s)\rangle \leq C_8. 
 \label{eqn:1.227-2}
\end{equation} 
We have also 
\[ \tilde s^{-1/2}\vert y_j(\tilde s)\vert\leq C_{10}, \quad \lim_{\tilde s\uparrow+\infty}\tilde s^{-1/2}b_j(\tilde s)=0 \] 
in (\ref{eqn:ext14}). 

Similarly to the remark after Lemma \ref{pro:1}, the singular part of $\tilde \zeta(dy,s)$, $s\in {\bf R}$, denoted by $\tilde \zeta_s(dy,s)$, is composed of a finite sum of delta fucntions.  By applying Lemma \ref{thm:ss02a} to the scaling back $\tilde A(dy,s)$ of $\zeta(dy,s)$, defined as in (\ref{eqn:c9}), the coefficient of each delta function of $\tilde \zeta_s(dy,s)$ must be less than or equal to $8\pi$.  These properties guarantee that the singular support of $\tilde \zeta(dy,s)$, denoted by ${\cal S}_s$, is composed of a finite number of collisionless accumulating points of $\{ \tilde s^{-1/2}y_j(\tilde s) \mid 1\leq j\leq m(\tilde s)\}$ defined for $-\log\tilde s=s+\tilde s_\ell$ as $\ell \rightarrow\infty$.  We may assume also that ${\cal S}_s$, $s\in {\bf Q}$, is composed of their converging points  by a diagonal argument. 

Therefore, we have $m(s)\in {\bf N}\cup\{ 0\}$ and $y^j_\infty(s)\in {\bf R}^2$, $1\leq j\leq m(\tilde s)$,  such that 
\[ \displaystyle{ {\cal S}_s=\{ y^j_\infty(s) \mid 1\leq j \leq m(s)\}}, \quad \tilde \zeta(\{ y^j_\infty(s)\}, s)=8\pi \] 
for $s\in {\bf R}$.  Furthermore, the density function $g=g(y,s)$ of the absolutely continuous part of $\tilde \zeta(dy,s)$ is provided with the properties 
\begin{eqnarray} 
& & 0\leq g=g(y,s) \leq C_{11}, \ \int_{{\bf R}^2}\vert y\vert^2g(y, s) \ dy \leq C_8, \ \Vert g(\cdot, s)\Vert_1\leq m(x_0) \nonumber\\ 
& & \Vert g(\cdot,s)\Vert_1\geq \delta, \ s\in {\bf Q}. 
 \label{eqn:38}
\end{eqnarray} 
It holds also that ${\cal S}_s\subset B_R$ for any $R>C_{10}$.  

Since $\tilde \zeta(dy,s)\in C_\ast(-\infty, +\infty; {\cal M}({\bf R}^2))$ the set $G=\bigcup_s({\bf R}^2\setminus {\cal S}_s)\times \{ s\}$ is open in ${\bf R}^2\times (-\infty, +\infty)$.  Furthermore, the above $g=g(y,s)$ is smooth in $G$ and it holds that 
\begin{eqnarray} 
& & g_s=\Delta g-\nabla\cdot g\nabla w \quad \mbox{in $G$} \nonumber\\ 
& & w(y,s)=\frac{\vert y\vert^2}{4}+4\sum_{j=1}^{m(s)}\log\frac{1}{\vert y-y_\infty^j(s)\vert}+\Gamma\ast g. 
 \label{eqn:1.229}
\end{eqnarray} 
Let 
\begin{equation} 
v(y,s)=\vert y\vert^2/4+(\Gamma\ast g)(y,s). 
 \label{eqn:v}
\end{equation} 
Henceforth, we put 
\begin{equation}
\overline{f}(r)=\frac{1}{2\pi}\int_0^{2\pi}f(re^{\imath \theta})d\theta 
 \label{eqn:41}
\end{equation} 
for $f=f(y)$, $y\in {\bf R}^2$, where $y=re^{\imath \theta}$ is the polar coordinate.  

We take $B_r=B(0,r)$, $r>R$, and $s_0\in {\bf R}$, to set 
\[ \displaystyle{ B^\varepsilon_r=B_r\setminus \bigcup_{j=1}^{m(s_0)}B(y_\infty^j(s_0), \varepsilon)}. \] 
By Lemma \ref{thm:ss02a}, again, any collision of the points in ${\cal S}_s$ does not occur as $s$ varies.  Therefore, making $0<\varepsilon\ll 1$, we obtain $\sharp \left(B(y^j_\infty(s_0),\varepsilon)\cap {\cal S}_s\right)\leq 1$ and 
\[ y_\infty^i(s)\in B(y^j_\infty(s_0), \varepsilon) \quad \Rightarrow \quad (y-y_\infty^i(s))\cdot \nu_y\geq 0, \ y\in \partial B(y^j_\infty(s_0), \varepsilon) \] 
for $1\leq j\leq m(s_0)$ and $\vert s-s_0\vert \ll 1$, where $\nu_y$ denotes the outer unit normal vector. 

Then, (\ref{eqn:1.229}) implies 
\begin{equation} 
\frac{d}{ds}\int_{B_r^\varepsilon}g(y,s) \ dy \leq \int_{\partial B_r^\varepsilon}g_r(y,s) \ d\sigma_y -\int_{\partial B_r^\varepsilon}(gv_r)(y,s) \ d\sigma_y
 \label{eqn:1.232-1}
\end{equation} 
for $\vert s-s_0\vert \ll 1$, where $d\sigma=d\sigma_y$ denotes the line element. We integrate (\ref{eqn:1.232-1}) in $t$, to convert it to an inequality on the difference quotient with the mesh $h>0$. Now, making $\varepsilon\downarrow 0$ and then $h\downarrow 0$, we obtain 
\begin{equation} 
\frac{d^+}{ds}\int_{B_r}g \ dy \leq \int_{\partial B_r}g_r \ d\sigma-\int_{\partial B_r}(gv_r) \ d\sigma =\int_{\partial B_r}g_r-\frac{r}{2}g -g(\Gamma\ast g)_r \ d\sigma, 
 \label{eqn:1.239+2}
\end{equation} 
where 
\[ \frac{d^+}{ds}A(s)=\limsup_{h\downarrow 0}\frac{1}{h}(A(s+h)-A(s)). \] 
Then, inequality (\ref{eqn:1.239+2}) is valid to any $r>R$ and $-\infty<s<+\infty$. 

From $\Vert g(\cdot,s)\Vert_1+\Vert g(\cdot,s)\Vert_\infty\leq C_{12}$, it follows that 
\[ \sup_s\Vert (\nabla \Gamma\ast g)(\cdot,s)\Vert_\infty<+\infty. \] 
Then, (\ref{eqn:1.239+2}) implies 
\begin{eqnarray} 
& & \frac{d^+}{ds}\int_{B_r}g \leq \int_{\partial B_r}g_r-\frac{r}{2}g+C_{13}g \ d\sigma \nonumber\\ 
& & \quad = \frac{d}{dr}(r\int_0^{2\pi}g \ d\theta)-\int_0^{2\pi}(1+\frac{r^2}{2})g \ d\theta+C_{13}r\int_0^{2\pi}g \ d\theta,  
 \label{eqn:1.240+2}
\end{eqnarray} 
which means 
\begin{equation} 
\frac{d^+}{ds}\int_0^rr\overline{g}dr\leq \frac{d}{dr}(r\overline{g})-(\frac{r^2}{2}-C_{13}r+1)\overline{g}, \quad r>R, \ -\infty<s<+\infty, 
 \label{eqn:1.241+2}
\end{equation} 
recalling (\ref{eqn:41}). Here we have 
\begin{equation} 
B(r,s)\equiv \int_0^rr\overline{g} \ dr\geq \delta-\frac{C_{14}}{r^2+1}, \quad r>R, \ s\in {\bf Q} 
 \label{eqn:1.243-3}
\end{equation}  
by (\ref{eqn:38}).  The strong maximum principle, on the other hand, implies also $B(r,s)>0$ for any $(r,s)$.  

Inequality (\ref{eqn:1.241+2}) means 
\[ \partial_s^+B\leq B_{rr}-a(r)B_r, \quad a(r)=\frac{r}{2}-C_{13}+\frac{1}{r}. \] 
Using $\displaystyle{ A(r)=\frac{r^2}{4}-C_{13}r+\log r}$, we have $a=A'$ and hence 
\[ \partial_s^+(e^{-A}B)\leq (e^{-A}B_r)_r, \quad r>R, \ -\infty<s<+\infty \] 
by $B_{rr}-a(r)B_r=e^A(e^{-A}B_r)_r$.  Now we take $r_1>R$ such that 
\[ a(r)\geq \frac{r}{4}+1, \quad \delta-\frac{C_{14}}{r^2+1}\geq \delta/2, \qquad r\geq r_1. \] 

Let $0\leq \varphi=\varphi(r)$ be a $C^1$ function on $[r_1,\infty)$, piecewise $C^2$, satisfying $\varphi(r_1)=0$, $\varphi(r)>0$ for $r>r_1$, and 
\begin{equation} 
\int_{r_1}^\infty e^{-A}\varphi \ dr<+\infty, \quad 
\lim_{r\uparrow+\infty}e^{-A}(\varphi+\varphi_r)=0. 
 \label{eqn:1.243+3}
\end{equation} 
Then we put $\varphi(r)=0$ for $r\in [0,r_1]$. Since $B\geq 0$, $B_r\geq 0$, $B(\infty,t)\leq C_{15}$, each $s\in {\bf R}$ admits $r_j\uparrow+\infty$ such that $B_r(r_j,s)\rightarrow 0$, which guarantees 
\begin{eqnarray} 
& & \frac{d^+}{ds}\int_0^\infty e^{-A}B\varphi \ dr\leq \int_0^\infty(e^{-A}B_r)_r\varphi \ dr=-\int_0^\infty e^{-A}B_r\varphi_r \ dr \nonumber\\ 
& & \quad \leq \int_0^\infty(e^{-A}\varphi_r)_rB \ dr =\int_0^\infty (\varphi_{rr}-a\varphi_r)e^{-A}B \ dr.  
 \label{eqn:1.243+2}
\end{eqnarray} 
We impose, furthermore, that the existence of $\mu>0$ such that 
\begin{equation} 
\varphi_{rr}-a\varphi_r\leq -\mu\varphi, \quad r\geq r_1 
 \label{eqn:1.243-2} 
\end{equation}
To assure all the above requirements to $\varphi(r)$, we take $0<c_1\ll 1$, for example, and put 
\[ \varphi(r)=\left\{ \begin{array}{ll} 
c_1r+c_2, & r\geq r_2 \\ 
\sin\beta(r-r_1), & r_1\leq r<r_2\equiv r_1+\frac{\pi}{4\beta} \end{array} \right. \] 
where 
\[ r_2=r_1+\frac{\pi}{4\sqrt{2}c_1}, \quad c_2=\frac{1}{\sqrt{2}}(1-\frac{\pi}{4})-r_1c_1, \quad \beta=\sqrt{2}c_1. \] 
Then we see $0\leq \varphi=\varphi(r)\in C^1[r_1,\infty)$, $\varphi(r_1)=0$, and (\ref{eqn:1.243+3}). Making $c_1\downarrow 0$, on the other hand, we obtain $r_2\uparrow+\infty$.  Therefore, (\ref{eqn:1.243-2}) arises for $0<\mu\ll 1$ by 
\[ \varphi_{rr}=\left\{ \begin{array}{ll} 
0, & r\geq r_2 \\ 
-\beta^2\varphi, & r_1\leq r<r_2 \end{array} \right. \] 
and 
\[ a(r)=\frac{r}{2}-C_{13}+\frac{1}{r}, \quad \varphi_r=c_1>0,  \qquad r\geq r_2. \] 

Since (\ref{eqn:1.243+2}) is obtained we have 
\[ \frac{d^+}{ds}\int_0^\infty e^{-A}B\varphi \ dr \leq -\mu \int_0^\infty e^{-A}B\varphi \ dr, \quad s\in {\bf R}. \] 
This means 
\[ \frac{d^+}{ds}\left\{ e^{\frac{\mu}{2} s}\int_0^\infty e^{-A}B\varphi \ dr\right\} < 0, \quad s\in {\bf R} \] 
and hence  
\[ \lim_{s\uparrow+\infty}\int_0^\infty e^{-A}B(\cdot,s)\varphi \ dr=0. \] 
We have, on the other hand, 
\[ \int_0^\infty e^{-A}B(\cdot,s)\varphi \ dr \geq \frac{\delta}{2}\int_0^\infty e^{-A}\varphi \ dr >0, \quad s\in {\bf Q} \] 
by (\ref{eqn:1.243-3}), a contradiction.  
\end{proof}

\begin{flushright}
Takashi Suzuki \\ 
Division of Mathematical Science \\ 
Department of Systems Innovation \\ 
Graduate School of Engineering Science \\ 
Osaka University \\ 
Machikaneyamacho 1-3 \\ 
Toyonakashi, 560-8531, Japan \\ 
suzuki@sigmath.es.osaka-u.ac.jp 
\end{flushright} 

\end{document}